OpenStudy (anonymous):

??? Which of the following is the simplified form of the quantity 4 y squared minus 1 over the quantity 2 y squared plus 3 y minus 2 ? 2y minus 1 over y minus 2 , with the restrictions y ≠ –1 over 2 and y ≠ –2 2y plus 1 over y plus 2 , with the restrictions y ≠ 1 over 2 and y ≠ 2 2y minus 1 over y minus 2, with the restriction y ≠ –1 over 2 and y ≠ 2 2y plus 1 over y plus 2, with the restriction y ≠ 1 over 2 and y ≠ –2

OpenStudy (swissgirl):

Lets write out the original equation as \( \large \frac{ 4y^2-1}{2y^2+3y-2}\)

OpenStudy (anonymous):

Okay,

OpenStudy (swissgirl):

So now we have to factor the numerator and the denomanator So lets start with the numerator \(4y^2-1\) which we know has the form \((a^2-b^2)=(a+b)(a-b)\) So in our case a^2=4y^2 so a=2y and b^2=1 so b=1 Therefore \(4y^2-1=(2y+1)(2y-1)\)

OpenStudy (anonymous):

Okai

OpenStudy (swissgirl):

Next we factor the denomanator \(2y^2+3y-2\) A=2 B=3 and C=-2 A*C=-4 So what 2 numbers when you multiply them it equals -4 and when you add them it equals 3 so the numbers are 4 and -1 4*-1=-4 4-1=3 So \( 2y^2+4y-y-2\) \(2y(y+2)-1(y+2)\) \( (2y-1)(y+2)\) So now we can rewrite our equation

OpenStudy (anonymous):

OK.

OpenStudy (swissgirl):

\(\huge \frac{ 4y^2-1}{2y^2+3y-2}=\frac{(2y+1)(2y-1)}{(2y-1)(y+2)}=\frac{(2y+1)}{(y+2)}\)

OpenStudy (swissgirl):

Now you gotta choose the right sentence

OpenStudy (anonymous):

Okay thank you.